# KeepKV: Eliminating Output Perturbation in KV Cache Compression for Efficient LLMs Inference

### v1

@@ 4.1 Experiment Settings
Tasks We evaluate KeepKV on datasets with standard and extended context lengths, covering question-answering, summarization, and synthetic tasks. Specifically, for question-answering, we utilize the COPA [34], MathQA [35], and OpenBookQA [36] tasks from the lm-eval-harness framework [37]. For summarization, we employ the XSUM[38] and CNN/DailyMail[39] tasks provided by the HELM framework. To assess performance on long-context tasks, we adopt LongBench[40], which effectively examines the algorithm’s compression capabilities across diverse subtasks, including single-document QA, multi-document QA, summarization, and synthetic tasks.
Models and baselines. Our evaluation is based on several representative LLMs, including OPT[41], Llama-2 [1], Llama-3 [2], and Mistral [3]. We compare our method against multiple baseline approaches: representative cache eviction methods such as Streaming [12], H2O [13] and PyramidInfer [18], and prominent cache merging methods including CaM [21] and D2O [23]. More detailed comparison results are provided in the Appendix.
Implementation. In our primary experiments, we set the merging threshold $T$ to $0.8$ and the exponential prediction coefficient $\beta=1.2$. For token selection and cache allocation, we follow the strategy recommended by PyramidInfer [18], which allocates fixed cache budgets, making it simple, and efficient. And it is sufficient to demonstrate the advantages of our algorithm. In contrast, D2O [23] applies dynamic allocation based on extra computation after prefill phase for each sequence. We implement KeepKV using the Hugging Face Transformers [42] and conduct all experiments on NVIDIA A100 80GB GPUs.
@@ 4.2 Accuracy on KV Cache Compression Ratios
Figure 5: Performance of KeepKV and other methods for LLama backbones on summarization datasets including XSUM and CNN/Daily Mail
@@ 4.3 Accuracy on Long-context Tasks
We evaluate KeepKV on the LongBench across Llama and Mistral model families . To thoroughly assess compression capabilities on long-context tasks, we set the KV cache budget ratio to $10\%$. The results indicate that KeepKV achieves performance closer to the full-cache baseline on most tasks, maintaining high generation quality despite limited cache availability. Notably, KeepKV significantly outperforms eviction-based methods. Furthermore, KeepKV also surpasses existing KV-cache merging methods, underscoring the effectiveness of our carefully designed merging strategy in enhancing output accuracy.
@@ 5 Conclusion
6 Acknowledgement
We are grateful to Chenhong He, Ruijie Miao, Yuhan Wu and Yanshu Wang from Peking University for their insightful discussions and helpful suggestions throughout the development of this research. We thank ByteDance Ltd. for providing technical support during the internship period.
@@ Appendix A Theoretical Analysis
Recently, many studies have analyzed KV cache compression strategies in LLM inference from a theoretical perspective [13, 15, 16, 18, 21, 23, 24, 43]. Overall, the primary objective of most existing works can be summarized as minimizing the impact of compression on the output. For instance, existing eviction-based methods and cache allocation strategies [13, 15, 14, 16, 17, 18, 19, 20, 31] all aim to maximize the retention of essential information within limited memory by evicting less important tokens or reducing cache allocation in non-critical heads and layers based on empirical observations of attention distributions. However, eviction inevitably leads to irreversible information loss, which has motivated the development of KV cache merging methods [21, 22, 23, 24]. Despite this, key challenges such as the selection of merging candidates and the assignment of merging weights remain largely unexplored, with a lack of systematic theoretical foundations. In this work, we introduce a novel perspective distinct from prior approaches. We formulate the problem as eliminating output perturbation and derive a novel merging method by analyzing the attention computation process. First, we introduce Electoral Votes mechanism, making the elimination of output perturbation feasible. Then, we derive a merging computation formula to eliminate perturbation at the current step. Finally, we extend this framework to multi-step generation, providing a theoretical guarantee for output perturbation and offering a reasonable explanation for mainstream similarity-based merging candidates selection methods.
Specifically, in Section A.1, we demonstrate the unavoidable output perturbation caused by KV cache eviction. In Section A.2, we discuss the attention sag issue in existing KV cache merging methods and provide a formal proof. In Section A.3, we present the derivation process of our KeepKV merging method. Finally, in Section A.4, we provide a theoretical guarantee for the output perturbation of KeepKV, including proofs for Theorem 3.5 and its associated lemma. The symbolic representation of attention computation process remains consistent with Section 3.
@@ A.1 Perturbation in KV Cache Eviction
Eviction methods discard KV pairs deemed unimportant. We denote the first generation step in the decoding phase as the $(L+1)$-th generation step, where $L$ represents the prompt length. And for a positive integer n, let $[n]:=\{1,2,...,n\}$. At $t$-th generation step, let $K_{e}=\{e_{1}e_{2},...,e_{m}\},m\in[t]$ denoted the index of to-be-evicted cache. Based on Equation 3, the output after eviction $(o^{\prime}_{t})$ is:
@@
However, existing studies [44] have shown that attention can be relatively dispersed in certain tasks, meaning that evicting even a small number of tokens can have a non-negligible impact. Furthermore, as the compression ratio increases, evicted tokens will account for a significant portion of the attention scores, exacerbating the degradation of generation quality.
@@ A.2 Attention Sag in KV Cache Merging
$k_{r}=w_{c}k_{c}+\sum_{j\in K_{e}}w_{j}k_{j},v_{r}=w_{c}v_{c}+\sum_{j\in K_{e}}w_{j}v_{j},\quad w_{c}+\sum_{j\in K_{e}}w_{j}=1$ (11)
Let $K^{\prime}_{e}=K_{e}\cup\{c\}$, representing the index of the original KVs before merging. For instance, the weight $w_{j}$ in D2O [23] is computed based on the cosine similarity between key vectors, whereas for KVmerger [24], it is calculated based on the Gaussian Kernel value. Formally, these are represented as follows:
$\ w_{j_{D2O}}=\frac{\exp(\cos\theta_{\mathbf{k_{j}},\mathbf{k_{c}}})}{\sum_{j\in K^{\prime}_{e}}\exp(\cos\theta_{\mathbf{k_{j}},\mathbf{k_{c}}})};\quad w_{j_{KVMerger}}=\frac{\exp(-\frac{||k_{j}-k_{c}||^{2}}{2\sigma^{2}})}{\sum_{j\in K^{\prime}_{e}}\exp(-\frac{||k_{j}-k_{c}||^{2}}{2\sigma^{2}})}.$ (12)
@@
Theorem A.1 (Formal version of Theorem 3.2).
@@
${A^{\prime}}_{r}^{t}=1-\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}}<1-\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}}{\sum_{i=1}^{t}s^{t}_{i}}=\sum_{i\in K^{\prime}_{e}}{A}_{i}^{t}$ (17)
@@
In the above expression, all vector coefficients are nonzero. Moreover, due to the high dimensionality and sparsity of the KV cache [24, 43], the vectors are almost linearly independent. In practical inference scenarios, it is nearly impossible for them to form a zero vector through linear combination. Consequently, we have: $\left\|o^{\prime}_{t}-o_{t}\right\|>0$. ∎
@@ A.3 KeepKV Merging Method
In Section 3.3.1, we introduced the concept of merging count via the Electoral Votes mechanism, aiming for a KV pair with vote count $p_{i}$ to be equivalent, in attention computation, to $p_{i}$ independent occurrences of this KV. Moreover, the vote count of the merged KV equals the sum of vote counts before merging. Formally, the outputs before $(o_{t})$ and after merging $(o^{\prime}_{t})$ can be expressed as follows:
$o_{t}=\frac{\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}}{\sum_{i=1}^{t}p_{i}s_{i}^{t}},\quad o^{\prime}_{t}=\frac{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r}}{\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}},p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}.$ (20)
Next, we demonstrate how our new merging approach in Section 3.3.1 can be derived naturally from the objective of eliminating output disturbances, which consequently serves as a direct proof for Theorem 3.3.
@@
$\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}=\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r},\quad\sum_{i=1}^{t}p_{i}s_{i}^{t}=\sum_{i=1,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}$ (21)
@@
$k_{r}=\frac{(\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}k_{i})\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}\ln{s_{i}^{t}}},\quad v_{r}=\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}v_{i}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}},\quad p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}$ (25)
Consequently, merging in this manner eliminates the output disturbance in the $t$-step, satisfying: $\left\|o^{\prime}_{t}-o_{t}\right\|=0$. By setting the merging candidates $K^{\prime}_{e}=\{e\}\cup\{c\}$, we obtain Theorem 3.3.

### v2

@@ 4.1 Experiment Settings
Tasks We evaluate KeepKV on datasets with standard and extended context lengths, covering question-answering, summarization, and synthetic tasks. Specifically, for question-answering, we utilize MathQA (amini2019mathqa), OpenBookQA (mihaylov2018openbookqa) and other tasks from the lm-eval-harness framework (lmevalharness). For summarization, we employ the XSUM(narayan2018xsum) and CNN/DailyMail(nallapati2016cnndm) tasks provided by the HELM framework. To assess performance on long-context tasks, we adopt LongBench(bai2024longbench), which effectively examines the algorithm’s compression capabilities across diverse subtasks, including single-document QA, multi-document QA and synthetic tasks.
Models and baselines. Our evaluation is based on several representative LLMs, including Llama-2 (touvron2023llama2), Llama-3 (grattafiori2024llama3), and Mistral (jiang2023mistral7b). We compare our method against multiple baseline approaches: representative cache eviction methods such as Streaming (xiao2024StreamingLLM), H2O (zhang2023h2o) and PyramidInfer (yang2024pyramidinfer), and prominent cache merging methods including CaM (zhang2024cam) and D2O (wan2024D2O). For all baselines, we follow the authors’ released implementations and hyperparameter settings in their papers.
Implementation. In our main experiments, we set the merging threshold $T$ to $0.8$. For token selection and cache allocation, we follow the strategy recommended by PyramidInfer (yang2024pyramidinfer), which allocates fixed cache budgets, making it simple and efficient. And it is sufficient to demonstrate the advantages of our algorithm. In contrast, D2O (wan2024D2O) applies dynamic allocation based on extra computation after prefill phase for each sequence. We implement KeepKV using the Hugging Face Transformers (wolf2020huggingfacestransformers) and experiments are by default conducted on NVIDIA A100 80GB GPUs.
Figure 4: Performance of KeepKV and other methods for LLama backbones on HELM and LM-Eval evaluations.
@@ 4.3 Accuracy on Long-context Tasks
We evaluate KeepKV on the LongBench across Llama and Mistral model families,including Llama-2-7B, Llama-2-13B, Llama-3-8B and Mistral-7B, as shown in Table 2. The evaluation tasks include Single-Document QA, Multi-Document QA, Summarization, Synthetic, and Code. The results indicate that KeepKV achieves performance closer to the full-cache baseline on most tasks, maintaining high generation quality despite limited cache availability. Notably, KeepKV significantly outperforms eviction-based methods, such as Local Window, StreamingLLM (xiao2024StreamingLLM), and H2O (zhang2023h2o). Furthermore, KeepKV also surpasses existing KV-cache merging methods, like CaM(zhang2024cam) and D2O(wan2024D2O), underscoring the effectiveness of our carefully designed merging strategy in enhancing output accuracy.
Methods Batch Size Throughput (tokens/s)
Full cache 2 116.54
H2O 8 317.33
D2O 8 214.8
KeepKV 8 255.99
Table 1: Throughput comparison of KeepKV and other methods (4k context, 20% compression ratio).
Methods Single-Doc QA Multi-Doc QA Summarization Synthetic Code
NrtvQA Qasper MF-en HotpotQA 2WikiMQA Musique TREC TriviaQA SAMSum PCount PRe Lcc RB-P
Llama-2-7B
Full Model 15.8 9.39 22.09 8.56 10.85 4.3 65.0 89.64 34.16 1.0 8.29 66.77 60.1
Local Window 2.22 9.29 1.83 5.14 7.18 1.02 17.5 4.07 3.17 1.5 2.58 16.31 15.35
StreamingLLM 11.81 5.18 19.26 7.07 10.48 3.71 55.5 87.31 31.84 1.5 4.29 63.79 56.07
H2O 16.54 7.57 20.61 7.68 9.28 4.09 64.0 87.98 33.62 1.34 9.14 65.34 58.49
CaM 11.79 5.1 19.12 7.26 10.48 3.64 56.0 87.31 31.85 1.5 4.29 63.66 55.98
D2O 16.04 6.54 19.48 8.14 10.12 4.62 63.5 88.39 34.1 1.39 7.54 65.8 59.44
Ours 17.32 7.48 22.2 8.51 9.72 4.65 60.5 88.87 33.2 2.23 8.45 65.9 56.36
Llama-2-13B
Full Model 12.64 8.61 19.82 9.1 10.98 5.8 69.5 87.04 41.89 2.0 6.03 67.08 57.53
Local 4.95 5.11 3.82 7.05 9.87 3.42 19.0 7.83 2.63 1.17 6.51 16.7 14.65
StreamingLLM 5.04 5.75 12.24 9.4 10.47 4.71 57.0 82.48 37.21 1.5 5.04 61.47 50.84
H2O 13.83 6.41 15.52 9.04 9.55 5.53 66.0 86.08 40.2 2.88 7.37 64.52 55.46
CaM 5.16 5.95 12.31 9.19 10.52 4.66 57.0 82.48 37.28 2.5 5.25 61.75 50.71
D2O 12.76 6.53 14.87 8.59 10.34 5.75 66.5 86.52 40.52 2.0 6.99 65.23 55.84
Ours 12.09 6.89 17.81 9.49 10.54 5.79 66.8 82.72 41.35 1.75 7.55 64.81 56.29
Llama-3-8B
Full Model 14.34 13.68 21.7 9.42 10.75 6.99 72 90.7 45.13 3.74 6.72 70.54 66.04
Local 2.14 6.69 5.17 6.16 5.0 2.42 34.25 30.5 10.66 2.36 2.0 28.91 24.52
StreamingLLM 10.43 7.84 13.85 9.18 10.44 5.47 61.0 90.37 44.35 2.6 10.5 68.49 63.94
H2O 13.73 10.02 17.2 9.31 10.62 6.42 63.3 90.44 45.02 3.29 7.56 68.95 63.84
CaM 10.43 7.83 13.89 9.11 10.37 5.47 61.0 90.37 44.31 3.16 10.5 68.59 64.04
D2O 13.5 8.86 17.21 9.16 10.52 6.35 65.5 90.52 44.64 3.44 5.8 68.49 64.84
Ours 12.76 10.63 18.57 9.37 10.72 6.53 64.5 90.33 45.2 3.54 7.16 69.05 65.68
Mistral-7B
Full Model 22.92 39.74 51.46 43.28 39.46 25.59 74.0 88.64 46.97 4.0 63.5 61.42 58.72
Local 16.89 16.92 21.11 23.33 22.49 10.23 58.5 81.29 36.3 2.1 7.71 41.1 47.88
StreamingLLM 16.76 17.28 21.41 24.16 22.54 10.72 60.3 82.21 37.43 2.14 7.67 51.19 47.94
H2O 18.06 16.75 22.28 24.77 21.68 8.86 61.0 83.03 30.34 2.15 5.76 56.5 49.88
CaM 16.46 17.26 21.4 25.66 22.54 10.72 59.17 82.21 37.33 2.14 7.67 51.01 47.89
D2O 18.58 15.92 21.71 26.41 21.68 9.07 61.5 83.12 39.5 2.18 7.3 57.51 50.59
Ours 18.16 17.95 22.93 26.56 23.18 9.42 62 83.47 39.7 2.19 7.26 58.9 50.71
Table 2: Performance evaluation of KeepKV on various models in LongBench benchmarks (20% compression ratio).
4.4 Throughput Analysis
Our experiments demonstrate that KeepKV significantly enhances the inference throughput of the model by efficiently compressing the KV Cache, as illustrated in Table 1. We conducted experiments on the Llama-2-7B model using an A100-80G GPU, with tasks derived from the LongBench evaluation framework. The experimental results indicate that various compression techniques improve throughput by reducing cache size and increasing batch size. Compared to the original full-cache method, KeepKV achieved over $2\times$ increase in throughput. It is noteworthy that, due to the additional computations, the throughput per request of merging methods is typically lower than that of classical eviction methods, such as H2O (zhang2023h2o). Nonetheless, KeepKV achieves higher throughput than the state-of-the-art (SOTA) merging-based algorithm, D2O (wan2024D2O). This advantage arises because D2O computes attention distribution variance for real-time cache allocation, whereas our method adopts a fixed cache allocation strategy to emphasizing the generalizability of KeepKV.
4.5 Ablation Study
To evaluate the generalizability of KeepKV, we conducted ablation experiments combining KeepKV with existing state-of-the-art eviction methods. Since KeepKV does not impose specific requirements on cache allocation or eviction/preservation strategies, it can be directly integrated with commonly used eviction/preservation policies. This only requires setting the merged parties as the eviction and preservation sets determined by their respective algorithms; it can also be applied with various cache allocation strategies, simply by modifying the cache configurations between layers and attention heads. As shown in Figure 5, we combined KeepKV with existing mainstream eviction methods, H2O(zhang2023h2o) and PyramidInfer(yang2024pyramidinfer), using the HELM evaluation framework. The results demonstrate that, with KeepKV incorporated, the methods outperform the original ones across all compression ratios. This proves that our algorithm is highly scalable and versatile, capable of being integrated with various eviction schemes to enhance their compression efficiency and generation quality.
(a) Combining with H2O.
(b) Combining with Pyramid.
Figure 5: Accuracy experiments combining KeepKV with existing eviction methods.
@@ 5 Conclusion
Acknowledgements
We are grateful to Chenhong He, Ruijie Miao, Yuhan Wu and Yanshu Wang from Peking University for their insightful discussions and helpful suggestions throughout the development of this research. We thank ByteDance Ltd. for providing technical support during the internship period. This work was supported by the National Key Research and Development Program of China under Grant No. 2024YFB2906603, by the Beijing Natural Science Foundation (Grant No. QY25123), and in part by the National Natural Science Foundation of China (NSFC) (624B2005).
@@ Appendix A Theoretical Analysis
Recently, many studies have analyzed KV cache compression strategies in LLM inference from a theoretical perspective (zhang2023h2o; liu2024scissorhands; li2024SnapKV; yang2024pyramidinfer; zhang2024cam; wan2024D2O; wang2024KVMerger; gu2025attentionsink_empiricalview). Overall, the primary objective of most existing works can be summarized as minimizing the impact of compression on the output. For instance, existing eviction-based methods and cache allocation strategies (zhang2023h2o; liu2024scissorhands; reid2024RoCo; li2024SnapKV; yang2024pyramidinfer; cai2024pyramidkv; feng2024adakv; fu2024headkv; xiao2024duoattention) all aim to maximize the retention of essential information within limited memory by evicting less important tokens or reducing cache allocation in non-critical heads and layers based on empirical observations of attention distributions. However, eviction inevitably leads to irreversible information loss, which has motivated the development of KV cache merging methods (zhang2024cam; wan2024D2O; wang2024KVMerger; nawrot2024dmc). Despite this, key challenges such as the selection of merging candidates and the assignment of merging weights remain largely unexplored, with a lack of systematic theoretical foundations. In this work, we introduce a novel perspective distinct from prior approaches. We formulate the problem as eliminating output perturbation and derive a novel merging method by analyzing the attention computation process. First, we introduce Electoral Votes mechanism, making the elimination of output perturbation feasible. Then, we derive a merging computation formula to eliminate perturbation at the current step. Finally, we extend this framework to multi-step generation, providing a theoretical guarantee for output perturbation and offering a reasonable explanation for mainstream similarity-based merging candidates selection methods.
Specifically, we first demonstrate the unavoidable output perturbation caused by KV cache eviction. Next, we discuss the attention sag issue in existing KV cache merging methods and provide a formal proof. Then, we present the derivation process of our KeepKV merging method. Finally, we provide a theoretical guarantee for the output perturbation of KeepKV, including proofs for the main theorem and its associated lemma. The symbolic representation of the attention computation process remains consistent with the one introduced in the methodology section.
@@ A.1 Perturbation in KV Cache Eviction
Eviction methods discard KV pairs deemed unimportant. We denote the first generation step in the decoding phase as the $(L+1)$-th generation step, where $L$ represents the prompt length. And for a positive integer n, let $n]:=\{1,2,...,n\}$. At $t$-th generation step, let $K_{e}=\{e_{1}e_{2},...,e_{m}\},m\in[t]$ denoted the index of to-be-evicted cache. Based on Equation 3, the output after eviction $(o^{\prime}_{t})$ is:
@@
However, existing studies (chen2024MagicPIG) have shown that attention can be relatively dispersed in certain tasks, meaning that evicting even a small number of tokens can have a non-negligible impact. Furthermore, as the compression ratio increases, evicted tokens will account for a significant portion of the attention scores, exacerbating the degradation of generation quality.
@@ A.2 Attention Sag in KV Cache Merging
$k_{r}$ $=w_{c}k_{c}+\sum_{j\in K_{e}}w_{j}k_{j},$
$v_{r}$ $=w_{c}v_{c}+\sum_{j\in K_{e}}w_{j}v_{j},$
s.t. $\quad w_{c}+\sum_{j\in K_{e}}w_{j}=1$ (11)
Let $K^{\prime}_{e}=K_{e}\cup\{c\}$, representing the index of the original KVs before merging. For instance, the weight $w_{j}$ in D2O (wan2024D2O) is computed based on the cosine similarity between key vectors, whereas for KVmerger (wang2024KVMerger), it is calculated based on the Gaussian Kernel value. Formally, these are represented as follows:
$w_{j_{\text{D2O}}}$ $=\frac{\exp(\cos\theta_{\mathbf{k_{j}},\mathbf{k_{c}}})}{\sum_{j\in K^{\prime}_{e}}\exp(\cos\theta_{\mathbf{k_{j}},\mathbf{k_{c}}})},$
[…]
@@
Theorem 7 (Formal version of Theorem 2).
@@
${A^{\prime}}_{r}^{t}$ $=1-\frac{\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}}{\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}+s^{t}_{r}}$
$<1-\frac{\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t}s^{t}_{i}}{\sum_{i=1}^{t}s^{t}_{i}}=\sum_{i\in K^{\prime}_{e}}{A}_{i}^{t}$ (17)
@@
In the above expression, all vector coefficients are nonzero. Moreover, due to the high dimensionality and sparsity of the KV cache (wang2024KVMerger; gu2025attentionsink_empiricalview), the vectors are almost linearly independent. In practical inference scenarios, it is impossible for them to form a zero vector through linear combination. Consequently, we have: $\left\|o^{\prime}_{t}-o_{t}\right\|>0$. ∎
@@ A.3 KeepKV Merging Method
In the main text, we introduced the concept of merging count via the Electoral Votes mechanism, aiming for a KV pair with vote count $p_{i}$ to be equivalent, in attention computation, to $p_{i}$ independent occurrences of this KV. Moreover, the vote count of the merged KV equals the sum of vote counts before merging. Formally, the outputs before $(o_{t})$ and after merging $(o^{\prime}_{t})$ can be expressed as follows:
$o_{t}$ $=\frac{\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}}{\sum_{i=1}^{t}p_{i}s_{i}^{t}},$
$o^{\prime}_{t}$ $=\frac{\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r}}{\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}},$
$p_{r}$ $=\sum_{i\in K^{\prime}_{e}}p_{i}.$ (20)
Next, we demonstrate how our new merging approach can be derived naturally from the objective of eliminating output disturbances, which consequently serves as a direct proof for Theorem 3.
@@
$\sum_{i=1}^{t}p_{i}s_{i}^{t}v_{i}$ $=\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}v_{i}+p_{r}s_{r}^{t}v_{r},$
$\sum_{i=1}^{t}p_{i}s_{i}^{t}$ $=\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t}p_{i}s_{i}^{t}+p_{r}s_{r}^{t}.$ (21)
@@
$k_{r}$ $=\frac{\left(\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}k_{i}\right)\ln{\left(\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}}{\sum_{i\in K^{\prime}_{e}}p_{i}}\right)}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}\ln{s_{i}^{t}}},$
$v_{r}$ $=\frac{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}v_{i}}{\sum_{i\in K^{\prime}_{e}}p_{i}s_{i}^{t}},\quad p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}.$ (25)
Consequently, merging in this manner eliminates the output disturbance in the $t$-step, satisfying: $\left\|o^{\prime}_{t}-o_{t}\right\|=0$. By setting the merging candidates $K^{\prime}_{e}=\{e\}\cup\{c\}$, we obtain Theorem 3.